OpenStudy (darkigloo):

If dy/dx=1/(y+x^2) and y(0)=2, find an equation for y in terms of x.

OpenStudy (darkigloo):

Sorry, I mean dy/dx = 1/(y+x^2y)

OpenStudy (darkigloo):

I know that I have to separate the variables. So (y+x^2y)dy=dx. I don't know what to do next.

OpenStudy (anonymous):

factor out y

OpenStudy (darkigloo):

is this right? ydy=dx/1+x^2

OpenStudy (anonymous):

yeah

OpenStudy (darkigloo):

ok now is this next step correct, after i do the integral? y^2/2 = ln|1+x^2|

OpenStudy (anonymous):

try arctan(x)

OpenStudy (darkigloo):

what???

OpenStudy (anonymous):

the integral of 1/(1+x^2) is arctan(x)

OpenStudy (darkigloo):

oh. why can't i use the rule int(du/u)=ln|u|+c?

OpenStudy (anonymous):

du = 2x dx, but you don't have an extra x on the numerator

OpenStudy (darkigloo):

ok so then i find c right? which is 2. and i plug it in and get \[y=\sqrt{2(arctanx+2)}\]